# Tree-Based versus Hybrid Graphical-Textual Model Editors: An Empirical Study of Testing Specifications

### v1

@@ Abstract.
Tree-based model editors and hybrid graphical-textual model editors have advantages and limitations when editing domain models. Data is displayed hierarchically in tree-based model editors, whereas hybrid graphical-textual model editors capture high-level domain concepts graphically and low-level domain details textually. We conducted an empirical user study with 22 participants, to evaluate the implicit assumption of system modellers that hybrid notations are superior, and to investigate the tradeoffs between tree-based and hybrid model editors. The results of the user study indicate that users largely prefer hybrid editors and are more confident with hybrid notations for understanding the meaning of conditions. Furthermore, we found that tree editors provide superior performance for analysing ordered lists of model elements, whereas activities requiring the comprehension or modelling of complex conditions are carried out faster through a hybrid editor.
Hybrid Notations, Model Editors, Fuzz Testing, Empirical Study, Domain-Specific Languages, Model-Driven Engineering, Robotics
copyright: acmlicensedjournalyear: 2024doi: XXXXXXX.XXXXXXXconference: ACM/IEEE 27th International Conference on Model Driven Engineering Languages and Systems; September 22–27, 2024; Linz, Austriaisbn: 978-1-4503-XXXX-X/18/06ccs: Software and its engineering Domain specific languagesccs: Software and its engineering Visual languagesccs: Software and its engineering Empirical software validation
@@ 1. Introduction
Throughout this paper, we use the term model editor rather than the term workbench, as a workbench covers a broader scope, incorporating a model editor in addition to other components such as menus, toolbars and perspectives (Burnette, 2005). For brevity, the term hybrid will be used instead of the term hybrid graphical-textual, and additionally, the term editor will refer to a model editor.
@@
Choosing between a tree-based and a hybrid editor for editing domain models is not a straightforward decision, as both have advantages and drawbacks. Tree editors display domain data hierarchically, in a structured manner, by exposing parent-child relationships. Diagrammatic editors contained within hybrid editors do not provide explicit structure, and they are not appropriate for displaying parent-child relationships of multiple levels. For instance, parent elements can be displayed as containers that incorporate their child elements. However, if the children have corresponding child elements, which in turn have child elements as well, displaying all these elements in a diagram can hinder comprehension, as they would effectively be displayed as a container that contains other containers, that in turn contain other containers incorporating a set of leaf child elements. Thus, in such cases, it may be preferred to use multiple diagrams, e.g., by instantiating new diagrams containing the child elements of a parent element. Consequently, the domain data is split across multiple diagrams, and users may face difficulties in being able to relate them. A common drawback of hybrid editors is the initial cost of learning the associated part-graphical and part-textual syntaxes. However, once learned, models may potentially be edited faster, considering also that typically fewer clicks are required due to the definition of textual expressions (Addazi et al., 2017).
@@
Acknowledging the importance of selecting the most suitable editor to support effective and efficient modelling, in this paper we investigate the tradeoffs between tree-based and hybrid editors. System modellers often take for granted the assumption that hybrid notations are superior, however, this claim is not yet supported by empirical evidence. To evaluate this implicit assumption, we conducted an empirical user study to analyse the editor type preferred by system modellers. We recruited 22 subjects, that are either students or professionals working in industry or academia, with various levels of knowledge and expertise in computer programming, modelling, and model-driven engineering (MDE). In the user study, the participants used a DSL that enables the definition of testing specifications of robotic systems in simulation. Each participant used the DSL via both a tree editor and a hybrid editor, to execute a set of tasks that are inspired from an industrial case study of a large European Commission research project named Secure and Safe Multi-Robot Systems (SESAME (SESAME, 2024)). The user study aims to evaluate the performance and preferences of users, and provide insights on the usage of hybrid editors and tree editors.
@@ 2. Background
Hybrid DSLs and their supporting editors can be developed using Graphite (Predoaia et al., 2023; Predoaia, 2023), a tool that facilitates the systematic engineering of hybrid editors via automatic code generation. Graphite (https://github.com/epsilonlabs/graphite) uses Sirius (Eclipse Foundation, 2024b) for the specification of the graphical part of the syntax, and Xtext (Eclipse Foundation, 2024c) for defining the textual part of the syntax. Alternatively, projectional editors, such as JetBrains MPS (JetBrains, 2024), can be employed for developing such hybrid editors. The research from (Scheidgen, 2008; Obeo and TypeFox, 2017; Cooper, 2018; Altran, 2022), which are based on hand-written code, introduce techniques for engineering such kind of hybrid editors. However, they do not address consistency enforcement between textual expressions and graphical model elements. Furthermore, except (Cooper, 2018), no techniques have been proposed for uniformly reporting errors. Other streams of work involve using blended modelling (Ciccozzi et al., 2019) and Capella (Eclipse Foundation, 2024a); however, they do not address hybrid DSLs, as justified in (Predoaia et al., 2023).
Testing Specification. Multi-robot systems (MRS) are distributed and interconnected robotic teams capable of collectively performing tasks that are beyond the competency of a single robot (Parker, 2013). Simulation-based testing (SBT) enables investigating the capacity of an MRS to operate dependably using a virtual environment that closely resembles the real deployment environment. Modern robotic frameworks enable realistic simulation of robots, with strong conformance to physical hardware making the software deployment to real robots easier (Timperley et al., 2018). Compared to field testing, simulation-based testing supports searching large design spaces of components and system configurations, using limited physical hardware resources, thus reducing significantly the overheads for detecting potential violations of safety requirements (Harbin et al., 2021; Gerasimou et al., 2019).
Fuzz testing is an automated software testing technique that involves providing unexpected, invalid or random data as input to a computer program, with the purpose of finding exceptions, crashes and potential memory leaks (Godefroid, 2020). Fuzz testing is considered among the most efficient approaches for finding security vulnerabilities in software systems (Liang et al., 2018). In the context of our work, fuzz testing is applied as a simulation-based testing technique, by injecting intentional errors or faulty messages into a robotic simulation to monitor its responses and measure its robustness.
@@ 3. Case Study
The use case is an industrial case study in gearbox manufacturing provided by our SESAME industrial partner, who manufactures industrial robots and factory automation systems. This use case incorporates mobile robots to assemble gearboxes, together with human operators surrounding the assembly cell who assist in the process. During testing, we aim to obtain evidence to validate the safety requirement that robots will not generate hazards to human workers at operation. The primary safety consideration is to ensure that robots do not enter regions that could cause a collision hazard to human operators placed around the perimeter of the assembly cell frame. Under normal circumstances, motion trajectories are pre-planned to ensure conformity to this safety requirement. However, faults in communication, sensing, or motor interlock actuation may lead to robots malfunctioning and entering these forbidden areas. The simulation scenario was supplied for testing by end users, and implemented using a proprietary simulation platform provided by a SESAME industrial partner to be used as a digital twin (Partners, 2023a).
@@
Figure 1 shows the layout of the assembly cell. The primary safety requirement is that the robot grippers are forbidden to enter static safety zones surrounding the cell, in order to prevent hazards to human workers. The safety zones are illustrated as blue cuboid regions (Z1, Z2, Z3) on the right-hand side of the external cell frame in Figure 1. If the safety zone surrounding the gripper of the right-hand robot (green cylinder labeled ZR3200) collides with these regions, it will trigger violations of the static safety requirement. An output metric is used to quantify violations of the principal safety requirement, by tracking the occurrence of collisions of the gripper safety zone with any of Z1, Z2 or Z3. Following a safety analysis, we have identified the communication from the controller to the robot arm joints as vulnerable to message disruptions. Therefore, the primary safety requirement will be tested by applying fuzzing, altering the target motion value for robot arm joints with targeted uniform random jitter around the intended point.
We employ a DSL to raise the level of abstraction and make testing specification more comprehensible to robotics engineers. Modelling testing specifications via a DSL tailored to this specific domain is highly desirable, enabling non-expert users to easily test robotics systems (Casalaro et al., 2022; Bruyninckx et al., 2013; Harbin et al., 2023). Participants of the user study will apply the DSL in the context of this case study, via a tree editor and a hybrid editor, to investigate the following research questions:
@@ 4. Fuzzing-Based Testing DSL
This section presents the fuzzing-based testing DSL developed in the context of the SESAME project, and applied to the case study from Section 3. Through the DSL, users can define testing specifications for an MRS simulation, which control how fuzzing is applied in MRS testing. We present next the abstract syntax and the concrete syntaxes of the DSL and the SBT methodology enabled by the DSL.
@@ 4.1. Abstract Syntax
The abstract syntax (i.e., the metamodel of the DSL) has been defined with the Eclipse Modelling Framework (EMF) through the Ecore metamodelling language. Figure 2 shows an excerpt from the metamodel of the DSL as a UML diagram. The core meta-classes are TestingSpace, TestCampaign and Test. A TestingSpace (highlighted) represents the root element of the domain and comprises permissible fuzzing operations (possibleFuzzingOperations) that bound the potential fuzzing space. The testing space also includes performance metrics, which are used to quantify the robotic system performance for safety violations. A TestCampaign specifies and constrains an individual experiment, referencing from the TestingSpace a subset of metrics to use in campaign evaluation, together with a subset of chosen fuzzing operations in includedOperations. The performedTests attribute in TestCampaign is automatically populated during the experiment with dynamically generated and executed Tests.
@@
The FuzzingOperation class is a specific fuzzing operation, i.e., an entity that represents a specific strategy for making runtime modifications to the MRS. The property variableToAffect selects the MRS variable to which fuzzing operations are applied. These variables will be automatically manipulated by the testing platform when fuzzing is active. Subclasses of FuzzingOperation represent distinct fuzzing semantics; e.g. RandomValueFromSetOperation enables randomised changes to structured parameters within a message.
A CampaignResultSet has a name and represents either final or partial intermediate testing results, as specified by the ResultSetStatus enumeration. The results are shown as references to notable tests under results. In an evolutionary experiment, the set of non-dominated tests in the final campaign result set constitute the Pareto front of optimal metrics obtained during the experiment. The inclusion of a Test in the non-dominated result sets shows that no other tests dominate it with respect to the set of metrics used.
Condition-based fuzzing provides the ability to activate and deactivate fuzzing based on custom scenario-specific events, incorporating complex logic and spatial or geometric properties. A condition is a logical expression built hierarchically from comparison and logical operators, incorporating state variables from the robotic system. The values of condition variables change dynamically during test execution, with every variable value supplied from a Metric of corresponding name. Activation of a condition-based fuzzing operation is controlled by its start and end conditions. When the start condition of a fuzzing operation evaluates to true, that fuzzing operation will be activated, and will manipulate simulation events. Once activated, the fuzzing operation remains active until its end condition evaluates to true.
Condition-based fuzzing experiments aims at discovering these conditions automatically, identifying high-level MRS events that lead to failures or safety violations. The primary role of a system tester will be to inspect, parse and understand these auto-generated conditions. However, during debugging and exploratory testing, users may modify existing conditions or create new from scratch, to verify the impact of these changes upon the behaviour of a test.
Figure 4. Tree Editor — A Testing Space
@@ 4.2. Concrete Syntaxes
The fuzzing-based testing DSL can be used through two concrete syntaxes: a graphical tree-based syntax and a hybrid syntax. The graphical syntax is provided by the built-in EMF tree-based editor, whereas the hybrid syntax has been developed using Graphite (Predoaia et al., 2023).
Tree editor. A subset of a testing space used in the case study is shown in Figure 4. The TestCampaign is selected in the editor, and its attributes (e.g., includedOperations and metrics) are displayed in the Properties View from Figure 6. The test campaign contains three executed tests and two campaign result sets. Recall that a CampaignResultSet represents a logical grouping of executed tests. The first test contains the RandomValueFromSetOperation under includedOperations in the TestCampaign, and also two metric instances: one for fuzzingOperationTimes and one for collisionOccurrence, as specified by the metrics attribute displayed in the Properties View from Figure 6. Figure 5 expands the RandomValueFromSetOperation named attackJointPosR3200ProductMove_Link0 and displays its child elements. This operation is activated when its start condition evaluates to true. The Condition element represents the root of the activation condition and it contains a CompositeCondition, in which the left side of the condition is a BasicCondition and the right side is another CompositeCondition.
@@
Hybrid editor. The Graphite-based concrete syntax is shown in Figures 7 and 8. A graphically defined testing space is illustrated in the diagram in Figure 7. The elements of the testing space are displayed as rectangles with various colours and symbols depending on each element type. Arrows highlight relationships between elements. To create new elements, users can drag and drop various symbols from the Palette on top of the diagram. The diagram displays the high-level elements of the testing space; however, users can view lower-level details by double-clicking on specific elements. By double-clicking on a Test contained in a TestCampaign, a new diagram is instantiated with the child elements of the Test. For instance, Figure 8 (left diagram) shows a hybrid editor that contains the child elements of the first test from the test campaign. The activation condition is specified with a textual expression in an embedded textual editor, with developer assistance features such as syntax highlighting and auto-completion. Also, users can double-click the TestCampaign from Figure 7 to view its campaign result sets in a newly instantiated diagram. Figure 8 (right diagram) illustrates the campaign result sets of the test campaign, ordered from first to last via arrows. By double-clicking a Test from the campaign result sets, a new diagram similar to the one from Figure 8 (left diagram) is instantiated, illustrating the details of the Test. We should note that the capabilities and behaviour of the hybrid editor, e.g., the instantiation of new diagrams via double-clicking, are shown to the participants of the user study via a video tutorial.

### v2

@@ Abstract.
Tree-based model editors and hybrid graphical-textual model editors have advantages and limitations when editing domain models. Data is displayed hierarchically in tree-based model editors, whereas hybrid graphical-textual model editors capture high-level domain concepts graphically and low-level domain details textually. We conducted an empirical user study with 22 participants to evaluate the implicit assumption of system modellers that hybrid notations are superior, and to investigate the tradeoffs between the default EMF-based tree model editor and a Sirius/Xtext-based hybrid model editor. The results of the user study indicate that users largely prefer the hybrid editor and are more confident with hybrid notations for understanding the meaning of conditions. Furthermore, we found that the tree editor provided superior performance for analysing ordered lists of model elements, whereas activities requiring the comprehension or modelling of complex conditions were carried out faster through the hybrid editor.
Hybrid Notations, Model Editors, Fuzz Testing, Empirical Study
journalyear: 2024copyright: acmlicensedconference: ACM/IEEE 27th International Conference on Model Driven Engineering Languages and Systems; September 22–27, 2024; Linz, Austriabooktitle: ACM/IEEE 27th International Conference on Model Driven Engineering Languages and Systems (MODELS ’24), September 22–27, 2024, Linz, Austriadoi: 10.1145/3640310.3674102isbn: 979-8-4007-0504-5/24/09ccs: Software and its engineering Domain specific languagesccs: Software and its engineering Visual languagesccs: Software and its engineering Empirical software validation
@@ 1. Introduction
Throughout this paper, we use the term model editor instead of the term workbench; a workbench covers a broader scope, including a model editor and other components like menus, toolbars and perspectives (Burnette, 2005). For brevity, we use the term hybrid instead of hybrid graphical-textual, and the term editor refers to a model editor.
@@
Choosing between a tree-based and a hybrid editor for editing domain models is not a straightforward decision, as both have advantages and drawbacks. Tree editors display domain data hierarchically and in a structured manner by exposing parent-child relationships. Diagrammatic editors contained within hybrid editors do not provide explicit structure, and they are not appropriate for displaying parent-child relationships of multiple levels. For instance, parent elements can be displayed as containers that incorporate their child elements. However, if the children have corresponding child elements, which in turn have child elements as well, displaying all these elements in a diagram can hinder comprehension, as they would effectively be displayed as a container that contains other containers, that in turn contain other containers incorporating a set of leaf child elements. Thus, in such cases, it may be preferred to use multiple diagrams, e.g., by instantiating new diagrams containing the child elements of a parent element. Consequently, the domain data is split across multiple diagrams, and users may face difficulties in being able to relate them. A common drawback of hybrid editors is the initial cost of learning the associated part-graphical and part-textual syntaxes. However, once learned, models may potentially be edited faster, considering also that typically fewer clicks are required due to the definition of textual expressions (Addazi et al., 2017).
@@
Acknowledging the importance of selecting the most suitable editor to support effective and efficient modelling, in this paper we investigate the tradeoffs between tree-based and hybrid editors. System modellers often take for granted the assumption that hybrid notations are superior, however, this claim is not yet supported by empirical evidence. To evaluate this implicit assumption, we conducted an empirical user study to analyse the editor type preferred by system modellers. We recruited 22 subjects, that are either students or professionals working in industry or academia, with various levels of knowledge and expertise in computer programming, modelling, and model-driven engineering (MDE). In the user study, the participants used a DSL that enables the definition of testing specifications of robotic systems in simulation. Each participant used the DSL via both a tree editor and a hybrid editor to execute a set of tasks that were inspired by an industrial case study of the large European Commission research project SESAME (Secure and Safe Multi-Robot Systems) (SESAME, 2024). The user study aims to evaluate the performance and preferences of users, and provide insights on the usage of hybrid editors and tree editors.
@@ 2. Background
Hybrid DSLs and their supporting editors can be developed using Graphite (Predoaia et al., 2023; Predoaia, 2023), a tool that facilitates the systematic engineering of hybrid editors via automatic code generation. Graphite (https://github.com/epsilonlabs/graphite) uses Sirius (Eclipse Foundation, 2024b) for the specification of the graphical part of the syntax, and Xtext (Eclipse Foundation, 2024c) for defining the textual part of the syntax. Alternatively, projectional editors, such as JetBrains MPS (JetBrains, 2024), can be employed for developing such hybrid editors. The research works from (Scheidgen, 2008; Obeo and TypeFox, 2017; Cooper, 2018; Altran, 2022), which are based on hand-written code, introduce techniques for engineering such kind of hybrid editors. However, they do not address consistency enforcement between textual expressions and graphical model elements. Furthermore, except (Cooper, 2018), no techniques have been proposed for uniformly reporting errors. Other streams of work involve using blended modelling (Ciccozzi et al., 2019) and Capella (Eclipse Foundation, 2024a); however, they do not address hybrid DSLs, as justified in (Predoaia et al., 2023).
Testing Specification. Multi-robot systems (MRS) are distributed and interconnected robotic teams capable of collectively performing tasks that are beyond the competency of a single robot (Parker, 2013). Simulation-based testing (SBT) enables investigating the capacity of an MRS to operate dependably using a virtual environment that closely resembles the real deployment environment. Modern robotic frameworks enable realistic simulation of robots, with strong conformance to physical hardware, making the software deployment to real robots easier (Timperley et al., 2018). Compared to field testing, simulation-based testing supports searching large design spaces of components and system configurations, using limited physical hardware resources, thus reducing significantly the overheads of detecting potential violations of safety requirements (Harbin et al., 2021; Gerasimou et al., 2019).
Fuzz testing is an automated software testing technique that involves providing unexpected, invalid or random data as input to a computer program, aiming to find exceptions, crashes and potential memory leaks (Godefroid, 2020). Fuzz testing is considered among the most efficient approaches for finding security vulnerabilities (Liang et al., 2018). In this work, fuzz testing is applied as a simulation-based testing technique to measure the robustness of a robotic mission.
@@ 3. Case Study
The use case is an industrial case study in gearbox manufacturing provided by our SESAME industrial partner, KUKA (KUKA, 2024), a company that manufactures industrial robots and factory automation systems. This use case incorporates mobile robots to assemble gearboxes, together with human operators surrounding the assembly cell who assist in the process. During testing, we aim to obtain evidence to validate the safety requirement that robots will not generate hazards to human workers during operation. The primary safety consideration is to ensure that robots do not enter regions that could cause a collision hazard to human operators placed around the perimeter of the assembly cell frame. Under normal circumstances, motion trajectories are pre-planned to ensure conformity to this safety requirement. However, faults in communication, sensing, or motor interlock actuation may lead to robots malfunctioning and entering these forbidden areas. The simulation scenario was supplied for testing by end users, and implemented using a proprietary simulation platform provided by our industrial partner and used as a digital twin (Partners, 2023a).
@@
Figure 1 shows the layout of the assembly cell. The primary safety requirement is that the robot grippers are forbidden from entering static safety zones surrounding the cell in order to prevent hazards to human workers. The safety zones are illustrated as blue cuboid regions (Z1, Z2, Z3) on the right-hand side of the external cell frame in Figure 1. If the safety zone surrounding the gripper of the right-hand robot (green cylinder labelled ZR3200) collides with these regions, it will trigger violations of the static safety requirement. An output metric is used to quantify violations of the principal safety requirement, by tracking the occurrence of collisions of the gripper safety zone with any of Z1, Z2 or Z3. Following a safety analysis, we have identified the communication from the controller to the robot arm joints as vulnerable to message disruptions. Therefore, the primary safety requirement will be tested by applying fuzzing, altering the target motion value for robot arm joints with targeted uniform random jitter around the intended point.
We employ a DSL to raise the level of abstraction and make testing specification more comprehensible to robotics engineers. Modelling testing specifications via a DSL tailored to this specific domain is highly desirable, enabling non-expert users to easily test robotics systems (Casalaro et al., 2022; Bruyninckx et al., 2013; Harbin et al., 2023; Wood et al., 2020). Participants of the user study will apply the DSL in the context of this case study, via a tree editor and a hybrid editor, to investigate the following research questions:
@@ 4. Fuzzing-Based Testing DSL
This section presents the fuzzing-based testing DSL developed in the context of the SESAME project, and applied to the case study from Section 3. Through the DSL, users can define testing specifications for an MRS simulation, which controls how fuzzing is applied during testing. Next, we present the abstract syntax and the concrete syntaxes of the DSL and the testing process.
@@ 4.1. Abstract Syntax
The abstract syntax (i.e., the metamodel of the DSL) has been defined with the Eclipse Modelling Framework (EMF) through the Ecore metamodelling language. Figure 2 shows an excerpt from the metamodel of the DSL as a class diagram. The core meta-classes are TestingSpace, TestCampaign and Test. A TestingSpace (highlighted) represents the root element of the domain and comprises permissible fuzzing operations (possibleFuzzingOperations) that bind the potential fuzzing space. The testing space also includes performance metrics, which are used to quantify the robotic system performance for safety violations. A TestCampaign specifies and constrains an individual experiment, referencing from the TestingSpace a subset of metrics to use in campaign evaluation, together with a subset of chosen fuzzing operations in includedOperations. The performedTests attribute in TestCampaign is automatically populated during the experiment with dynamically generated and executed Tests.
@@
The FuzzingOperation class is a specific fuzzing operation, i.e., an entity that represents a specific strategy for making runtime modifications to the MRS. The property variableToAffect selects the MRS variable to which fuzzing operations are applied. These variables will be automatically manipulated by the testing platform when fuzzing is active. Subclasses of FuzzingOperation represent distinct fuzzing semantics; e.g., RandomValueFromSetOperation enables randomised changes to structured parameters within a message.
A CampaignResultSet has a name and represents either final or partial intermediate testing results, as specified by the ResultSetStatus enumeration. The results are shown as references to notable tests under results. In an evolutionary experiment, the set of non-dominated tests in the final campaign result set constitutes the Pareto front of optimal metrics obtained during the experiment. The inclusion of a Test in the non-dominated result sets shows that no other tests dominate it with respect to the set of metrics used.
Condition-based fuzzing allows users to activate and deactivate fuzzing based on custom scenario-specific events, incorporating complex logic and spatial or geometric properties. A condition is a logical expression built hierarchically from comparison and logical operators, incorporating state variables from the robotic system. The values of condition variables change dynamically during test execution, with every variable value supplied from a Metric of the corresponding name. Activation of a condition-based fuzzing operation is controlled by its start and end conditions. When the start condition of a fuzzing operation holds, that fuzzing operation will be activated, and will manipulate simulation events. Once activated, the fuzzing operation remains active until its end condition holds.
Condition-based fuzzing experiments aim to discover these conditions automatically, identifying high-level MRS events that lead to failures or safety violations. The primary role of a system tester will be to inspect, parse and understand these auto-generated conditions. However, during debugging and exploratory testing, users may modify existing conditions or create new ones from scratch, to verify the impact of these changes on the behaviour of a test.
Figure 4. Tree Editor — Testing Space
@@ 4.2. Concrete Syntaxes
The fuzzing-based testing DSL can be used through two concrete syntaxes: a graphical tree-based syntax and a hybrid syntax. The graphical syntax is provided by the default EMF-generated tree editor, whereas the hybrid syntax was developed using Graphite (Predoaia et al., 2023).
Tree editor. A subset of a testing space used in the case study is shown in Figure 4. The TestCampaign is selected in the editor, and its attributes (e.g., includedOperations and metrics) are displayed in the Properties View from Figure 6. The test campaign contains three executed tests and two campaign result sets. Recall that a CampaignResultSet represents a logical grouping of executed tests. The first test contains the RandomValueFromSetOperation under includedOperations in the TestCampaign, and also two metric instances: one for fuzzingOperationTimes and one for collisionOccurrence, as specified by the metrics attribute displayed in the Properties View from Figure 6. Figure 5 expands the RandomValueFromSetOperation named attackJointPosR3200ProductMove_Link0 and displays its child elements. This operation is activated when its start condition holds. The Condition element is the root of the activation condition and it has a CompositeCondition, in which the left side is a BasicCondition and the right side is another CompositeCondition.
@@
Hybrid editor. The Graphite-based concrete syntax is shown in Figures 7 and 8. A graphically defined testing space is illustrated in the diagram in Figure 7. The elements of the testing space are displayed as rectangles with various colours and symbols depending on each element type. Arrows highlight relationships between elements. To create new elements, users can drag and drop various symbols from the Palette on top of the diagram. The diagram displays the high-level elements of the testing space; however, users can view lower-level details by double-clicking on specific elements. By double-clicking on a Test contained in a TestCampaign, a new diagram is instantiated with the child elements of the Test. For instance, Figure 8 (left diagram) shows a hybrid editor that contains the child elements of the first test from the test campaign. The activation condition is specified with a textual expression in an embedded textual editor, with developer assistance features such as syntax highlighting and auto-completion. Users can also double-click the TestCampaign from Figure 7 to view its campaign result sets in a newly instantiated diagram. Figure 8 (right diagram) illustrates the campaign result sets of the test campaign, ordered from first to last via arrows. By double-clicking a Test from the campaign result sets, a new diagram similar to the one from Figure 8 (left diagram) is instantiated, illustrating the details of the Test. We should note that the capabilities and behaviour of the hybrid editor, e.g., the instantiation of new diagrams via double-clicking, are shown to the participants of the user study via a video tutorial.